### v1

# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

Abstract.
Recently, a distributed middleware application called contract automata runtime environment (CARE) has been introduced to realise service applications specified using a dialect of finite-state automata. In this paper, we detail the formal modelling, verification and testing of CARE. We provide a formalisation as a network of stochastic timed automata. The model is verified against the desired properties with the tool Uppaal, utilising exhaustive and statistical model checking techniques. Abstract tests are generated from the Uppaal models that are concretised for testing CARE. This research emphasises the advantages of employing formal modelling, verification and testing processes to enhance the dependability of an open-source distributed application. We discuss the methodology used for modelling the application and generating concrete tests from the abstract model, addressing the issues that have been identified and fixed.
1. Introduction
Behavioural contracts [3] have been introduced to formally describe the interactions among services, to enable to reason formally about well-behaving properties of their composition. Examples of properties are agreement among the parties or reachability of target states.
Contract automata are a dialect of finite state automata used to specify behavioural contracts formally in terms of offers and requests [17]. A composition of contracts is in agreement when all requests are matched by corresponding offers of other contracts. A composition can be refined to one in agreement using the orchestration synthesis algorithm [13, 16], a variation of the synthesis algorithm from supervisory control theory [44]. The Contract Automata Runtime Environment (CARE[11] provides a middleware to coordinate the services implementing contracts. In CARE, each transition of the orchestration automaton is executed by a series of interactions among the orchestrator and the CARE services, implemented using Java TCP/IP sockets. These interactions may vary according to the specific configuration chosen among those provided by CARE. In [11] the algorithms implemented in CARE are proved to enforce the adherence of each contract specification to its CARE implementation (basically, the control flow of the application follows the synthesised orchestration automaton).
In this paper, we describe the modelling, verification and testing of the low-level interactions among CARE services and the orchestrator. This aspect is no less important, as witnessed by known cases of algorithms proved to be correct (e.g., the Byzantine distributed consensus [35]) and whose low-level communication implementations were found to have issues, e.g. deadlocks [45]. We verify several properties, including the absence of deadlocks, absence of undelivered messages, and reachability of target states.
The formal model is a network of stochastic timed automata as accepted by the Uppaal toolbox. Different variants of the formal model are proposed. The model undergoes verification against the desired properties using Uppaal, employing both exhaustive model checking, statistical model checking and model-based testing. This combination allows for thorough analysis of the model’s behavior and ensures scalability when dealing with systems that possess a large state-space.
This work is an extension of the conference paper [7]. The extension includes the addition of more figures and tables, as well as further explanations throughout the paper. Testing, which was only briefly mentioned in [7], is now thoroughly described. We cover both the generation of abstract tests from the Uppaal model and the process of concretising these tests in the real application.
The benefits of modelling and verifying CARE are: (i) increasing the confidence in the reliability of CARE due to the formal verification through model checking and model-based testing, (ii) quantitative evaluation of measures of interest of CARE obtained through statistical model checking, (iii) improved documentation thanks to the graphical and animatable state machines endowed with precise semantics.
Finally, this paper tackles the challenge of providing a full-fledged model-based development and formal methods approach [26]. The final application has been graphically modelled at an abstract level, formally verified and tested using the formal model. The criteria followed for abstracting away irrelevant details are discussed together with the issues that have been found and fixed thanks to the formal modelling, verification and testing.
All models, formulas, and logs are publicly available in [6], together with traceability and model-based testing information connecting the model to the source code.
Outline We start by discussing the related work in Section 2. In Section 3, we provide the background on contract automata, their tool support, Uppaal, and the contract automata runtime environment. The methodology used for modelling and analysing CARE is described in Section 4. Section 5 contains the description of the models, whilst the verification is described in Section 6. Model-based testing is addressed in Section 7. Finally, the conclusion and future work are presented in Section 8.
2. Related work
Several applications of Uppaal to various case studies are available in the literature, including land transport [15], maritime transport [47], medical systems [38], and autonomous agents path planning [28]. These case studies, along with the present paper, adopt a model-based approach, wherein partial representations of the applications are created using models.
A recent survey conducted on formal methods [25] reveals that numerous publications lack a direct correlation between the concrete implementation and its abstract model. This holds especially when the systems are only envisioned, yet to be realised, such as for example the ERTMS L3 railway signalling system [4, 14]. This is also the case for industrial systems whose implementations are non-disclosed [31]. In all these cases, determining the accuracy of the model in relation to the actual system and assessing the appropriateness of the chosen level of abstraction becomes challenging. Furthermore, measuring the influence of the formal modelling and verification phase on the analyzed system poses difficulties.
Indeed, a recent survey on formal methods in the transport domain [25] identifies the need “to lower the degree of abstraction, by applying formal methods and tools to development phases that are closer to software development”. This is also confirmed by a recent survey on formal methods among 130 high-profile experts [26], where nearly half of them believe that the most likely future users of formal methods are “a small number of skilled experts” as long as “skills like modelling, specification, abstraction are involved”.
In contrast to the previously mentioned literature, in this paper we present a bottom-up formal analysis of an established open-source system that has already been developed [11]. The availability of the source code further enables us to establish a connection between the abstract formal model and the actual source code. This capability facilitates the precise identification of specific aspects of the real system that have been abstracted in the formal model. Additionally, it allows us to validate the appropriateness of the chosen level of abstraction.
Recently, the offline model-based testing functionality of Uppaal, called Yggdrasil [30], has been used in [33] for testing an industrial real-time automotive turn indicator system, and in [29] for testing EIP-1559 Ethereum smart contracts. The usage of the Uppaal suite, and more specifically Yggdrasil, in three industrial case studies (i.e., medical devices, an automotive system and a pump manufacturing) is reviewed in [36]. Many of the above applications of model-based testing using Uppaal concern industrial, non-disclosed case studies. Therefore, only the models and abstract test cases are reported, whilst concrete test cases and source code have not been made public. In comparison, CARE is an open-source application. The concrete tests generated from the Uppaal models discussed in this paper, as well as the system under testing, are available at [6]. To the best of our knowledge, there are no other non-trivial open-source applications for which their Uppaal formal model is openly accessible and directly connected to the source code through traceability and model-based testing. This contribution assists in linking formal methods, particularly Uppaal, to the software development process.
The contract automata approach is closer to [34], where behavioural types are expressed as finite state automata of Mungo, called typestates [48, 1]. Similarly to CARE, in Mungo finite state automata are used as behaviour assigned to Java classes (one automaton per class), with transition labels corresponding to methods of the classes. A tool to translate typestates into automata was presented in [49]. CATApp is a graphical front-end tool for designing contract automata [5]. A tool similar to Mungo is JaTyC (Java Typestate Checker) [2].
In the literature there are various attempts to model applications that, similarly to CARE, communicates through TCP/IP sockets, also using Uppaal [50, 24, 46, 51]. In [22] the verification of the transport layer of a TCP/IP protocol stack implementation in C is performed with a translation into SPARK, where contracts are modelling the corresponding behaviour. A method to obtain SPIN formal models from TCP/IP socket applications implemented in C is presented in [21] where the C application is embedded into a PROMELA model equipped with a formalisation of the underlying TCP/IP socket API. Similarly, we provide a model of the underlying socket communications but we also modelled aspects of the application that are targeted by our analysis whilst unnecessary details are abstracted away. Another approach [39] presents an automatic translation from MAUDE specifications of distributed applications to distributed MAUDE implementations using TCP/IP sockets, where the implementations are formally proven to be in correspondence with the specifications. Our approach is bottom-up and is based on Uppaal modelling of an already realised application.
As reported in [20], the question of qualification and validation of formal methods tools is “absolutely crucial”. Other tools for behavioural contracts are present in the literature [27, 41, 42], but differently from CARE, many of these implementations have not undergone a process of formal verification.
3. Background
In this section, we will provide background on contract automata, their software support, and the Uppaal statistical model checker. The focus of this paper is on the formal analysis of the runtime environment of contract automata. While we offer a concise overview of contract automata to enhance comprehension of their runtime environment, they are not the focus of our formal analysis.
3.1. Contract Automata
Contract automata are a dialect of finite-state automata modelling services that exchange offers and requests. A contract automaton models either a single service or a composition of interacting services. Labels of transitions of contract automata are vectors of atomic elements called actions. Similarly, states of contract automata are vectors of atomic elements called basic states. The length of the vectors is equal to the number of services in the automaton (this number is called rank). A request (resp., offer) action is prefixed by ? (resp., !). The idle action is denoted by -. Labels are constrained to be one out of three types. In a request (resp., offer) label a service performs a request (resp., offer) action and all other services are idle. In a match label one service performs a request action, another service performs a matching offer action, and all other services are idle. For example, the contract automaton in Figure 8 right has rank 2 and the label [?quit,!quit] is a match where the request action ?quit is matched by the offer action !quit. Note the difference between a request label, e.g. [?coffee, -], and a request action, e.g. ?coffee.
In a composition of contracts various properties can be analysed [17]. For example, the property of agreement requires to match all request actions, whereas offer actions can remain unmatched. The synthesis of the orchestration in agreement produces a sub-automaton of the composition where all services can match their requests with corresponding offers to reach a final state. Thus, in the orchestration in agreement labels of transitions are only matches or offers [16]. The contract automaton in Figure 8 right is an orchestration in agreement.
Contract automata and their functionalities are implemented in a software artefact, called Contract Automata Library (CATLib[9]. With CATLib it is possible to specify, compose, and synthesise specifications given as contract automata. Indeed, CARE uses the facilities offered by CATLib for, among other things, composing contracts and synthesising orchestrations. This software artefact is a by-product of our scientific research on behavioural contracts and implements results that have previously been formally specified in several publications (cf., e.g., [13, 16, 17]). Scalability features offered by CATLib include a bounded on-the-fly state-space generation optimised with pruning of redundant transitions and parallel streams computations [10, 12]. The software is open source [9], it has been developed using principles of model-based software engineering [8] and it has been extensively validated using various testing and analysis tools to increase the confidence on the reliability of the library [9].
3.2. Uppaal
Uppaal SMC [23] is a variant of Uppaal [19], which is a well-known toolbox for the verification of real-time systems. Uppaal models are stochastic timed automata, in which non-determinism is replaced with probabilistic choices and time delays with probability distributions (uniform for bounded time and exponential for unbounded time). These automata may communicate via broadcast channels and shared variables. In this paper, we will use both exhaustive and statistical model checking. Statistical Model Checking (SMC) [37] involves running a sufficient number of (probabilistic) simulations of a system model to obtain statistical evidence (with a predefined level of statistical confidence) of the quantitative properties to be checked. Monte Carlo estimation with Chernoff-Hoeffding bound executes $N=\lceil(ln(2)-ln(\alpha))/(2\varepsilon^{2})\rceil$ simulations $\rho_{i}$, $i\!\in\!1...N$, to provide the interval $[p^{\prime}-\varepsilon,p^{\prime}+\varepsilon]$ with confidence $1-\alpha$, where $p^{\prime}=(\#\{\rho_{i}\!\mid\!\rho_{i}\!\models\!\varphi\})/N$, i.e., $\mathtt{Pr}(|p^{\prime}-p|\leq\varepsilon)\geq 1-\alpha$ where $p$ is the unknown value of the formula $\varphi$ being estimated statistically and $\varepsilon$ and $\alpha$ are the user-defined precision and confidence, respectively. Crucially, the number of simulations used to estimate a formula is independent of the model’s size and depends only on the parameters $\alpha$ and $\varepsilon$. In practice the number of simulations required by Uppaal to reach a specific confidence level is optimized and is thus lower than the above theoretical bound. Uppaal supports template automata used to instantiate different copies (in different experiments) of the same automaton, distinguishable by their parameters. The selection of Uppaal as the chosen tool is influenced by several factors. These include its extensive adoption by the community, expertise of the author, usability, primitive support for real-time and stochastic modeling, probabilistic and non-deterministic choices, and capabilities for statistical model checking, simulation, and model-based testing.
The off-line test case generator of Uppaal is called Yggdrasil, which is now integrated into the Uppaal GUI as a tab. Starting from the Uppaal models, a suite of test cases is created, aiming to cover all syntactic transitions in the model (i.e., edge coverage). The Uppaal model can be decorated with test code, which can be generated from the execution of a transition or at the entry to or exit from a location. Any language can be used for the test code. An abstract test case is generated from a trace of execution of the Uppaal model. During the execution of the trace, the test code of the model is dumped into a text file, forming the abstract test case. When the system is composed of a set of automata, it is possible to generate the test code from a single automaton or from the whole system. In this latter case, the single generated abstract test case contains the interleaved test code from all automata. The traces of execution can be generated as witnesses of some reachability property defined using the Uppaal syntax. Random traces can also be executed to improve test coverage. Parameters to be set include the depth of the trace, the type of search (i.e., breadth, depth), and options for the trace to be generated (i.e., fastest, shortest). The generated abstract test cases need to be decorated and filled with the missing information from the system under test to become concrete test cases.
3.3. CARE
The Contract Automata Runtime Environment (CARE[11] provides facilities for pairing the contract automata specifications with actual implementations of service-based applications. Note that our purpose is not to formally analyse the applications created using CARE, but to formally analyse CARE itself. The formal verification of applications developed using CARE is a consequence of the reliance of CARE on the formal guarantees provided by contract automata [11], along with the formal verification of CARE as an application. This paper addresses the latter aspect. This software is organised into classes for the orchestrated services (cf. Figure 1) and classes for the orchestrator.The two core abstract Java classes are the RunnableOrchestration and the RunnableOrchestratedContract.
Figure 1. The class diagram for the orchestrator; the methods of the derived classes are visible in their super-class/interface as abstract methods (in italic)
Services
In Figure 1, RunnableOrchestratedContract is an abstract class that implements an executable wrapper responsible for pairing its contract automaton specification (instance variable contract storing a contract automaton) with an implementation provided as a Java class (instance variable service implementing the service), where each action of the automaton contract is in correspondence with a method of the service class. The types of the methods of a matching offer and request must also be in correspondence. Each time a new orchestration involving the service is initiated, the RunnableOrchestratedContract creates a new service. This service remains in a waiting state to receive commands from the orchestrator, which then triggers the execution of the corresponding methods. In case the orchestrator requires to perform an action not prescribed by its service contract, then a ContractViolationException will be raised by the service. As demonstrated in [11], this exception will never be raised if the orchestration in agreement is synthesised using the algorithms of contract automata.
The realisation of an orchestration is abstracted away in contract automata. Crucially, offers and requests of contracts are an abstraction of low-level messages sent between the services and the orchestrator to realise them. CARE exploits the abstractions provided by Java to allow its specialisation according to different implementation choices, using abstractions of object-oriented design, as showed in Figure 1 and Figure 2.
Two aspects to implement are choices and termination (through the abstract method choice). CARE is equipped with default implementations, but can be extended (by implementing the relative interfaces and abstract methods) to include other options, other than the default ones. Currently, a so-called ‘dictatorial’ choice (i.e., an internal choice of the orchestrator, external for the services) and a so-called ‘majoritarian’ choice (services vote and the majority wins) are two implemented options. MajoritarianChoiceRunnableOrchestratedContract and DictatorialChoiceRunnableOrchestratedContract are the two classes specialising RunnableOrchestratedContract according to how the choice is handled and implementing the abstract methods. CARE also provides default implementations for the low-level message exchanges. Currently, the two available options are the ‘centralised’ action, where the orchestrator acts as a proxy, and the ‘distributed’ action, where two services matching their actions directly interact with each other once the orchestrator has made them aware of a matching partner and its address/port. Accordingly, each RunnableOrchestratedContract has an OrchestratedAction (instance variable act) used to implement the corresponding actions that can be either distributed or centralised according to the current implementation.
Figure 2. The class diagram for the orchestrated services; the methods of the derived classes are visible in their super-class/interface as abstract methods (in italic)
Orchestrator
The class diagram of the orchestrator side in Figure 2. The abstract class RunnableOrchestration implements a special service that reads the synthesised orchestration (stored in the instance variable contract) and orchestrates the services (the instances of RunnableOrchestratedContract) to realise the overall application. The instance variables port and addresses contain the corresponding addresses and ports of the services involved in the orchestration. Similarly to the case of the orchestrated contract, also the orchestrator is specialised according to either a dictatorial or a majoritarian implementation of the abstract method choice. Moreover, an OrchestratorAction instance variable is used to implement each action of the orchestration, either centralised or distributed, thus matching the corresponding actions of the orchestrated services.
4. Methodology
In this section, we describe both the methodology used for modeling the communications, abstracting away irrelevant details, and the adequacy of the model to the actual implementation.
4.1. Modelling TCP/IP sockets communication
Java TCP/IP sockets communications are asynchronous with FIFO buffers [40]. In Uppaal, the interactions are via channel synchronisation and global variables. Thus, TCP/IP sockets are solely employed in the actual implementation and are not utilized by the automata of the model. Instead, in the model each party communicates with the partner using two global arrays (one for sending and one for receiving, respectively). Figure 3 shows the code used in the Uppaal models (see Section 5) for modelling the TCP/IP socket communications.
The two global (FIFO) arrays in the model are named orc2services[N][queueSize] and services2orc[N][queueSize] (see Figure 3 left). These are used to model the TCP/IP sockets buffers, where N is the number of services in the model and queueSize is the size of the queue of the components. These arrays are only modified by functions for enqueueing and dequeuing messages.
Figure 3. The Uppaal code used to model the TCP/IP socket communications
The local declarations of the two automata contain methods for sending and receiving from the partners and checking their queues of messages. Both automata declare a method enqueue for sending to the partner a message. Messages are modeled as global integer constants, depicted in Figure 3 left. Indeed, in this model the actual payload of each communication is abstracted away The enqueue method of the services only takes as parameter the signal ide_sig to send (see Figure 3 center). The identifier of the service id is only needed on the orchestrator side to identify the partner (there is only one orchestrator thus no identifier is needed for it, see Figure 3 right). Similarly, both automata have a method dequeue for consuming messages from their respective arrays.
The default mode for Java TCP/IP sockets is blocking [32], meaning that the sender blocks when the buffer of the receiver is full, and waits until there is enough space to proceed. Accordingly, a transition having a send in its effect will check in its guard whether there is enough space left in the array of the partner by calling either the method available (returning the space left) or isFull. Similarly, in Java TCP/IP sockets the read operation is blocking. Accordingly, before reading it is always checked whether the array is not empty with the method !isEmpty. When the array is empty the automaton blocks until a message is received.
Source locations of sending and receiving transitions are neither committed nor urgent. In fact, an enabled committed transition (i.e., whose source location is committed, denoted with C) must be executed before any other non-committed transition in the network. Instead, an urgent transition must be executed without any delay. If a send or receiving transition were to be either committed or urgent, it would introduce the possibility of false positive deadlocks. This scenario arises when, for example, the receiver has a full buffer (i.e., array) and is prepared to free it, but the sender transition is enabled and committed. Similarly, this false positive can occur when the buffer is full, the send transition is urgent but the receive operation is not urgent, or vice versa, when the buffer is empty, the read transition is urgent, and the send operation is not urgent.
Hence, the operations of writing to and reading from a buffer are represented using stochastic delays, specifically following an exponential distribution. Two rates are employed to capture the delay associated with reading and writing. These exponential delays are present in all non-committed and non-final locations of both automata in Figure 6 and Figure 7. However, the presence of unbounded delays introduces scenarios where the receiver (resp., sender) may wait indefinitely without executing its read (resp., write) transition, even when it is enabled. These scenarios would invalidate the exhaustive model checking of reachability properties that are satisfied by the actual system, leading to false positives. In the real implementation, Java TCP/IP sockets offer a timeout mechanism wherein an exception is thrown if no message exchange occurs within a specified time frame. All sockets used by CARE have this timeout. Consequently, a dedicated automaton called SocketTimeout is replicated for each service to model the timeout operation (see Figure 8 left). Each send or receive operation in every socket resets the SocketTimeout clock c. If no reset operation is received within a certain duration (variable timeout), SocketTimeout enters a location called Timeout and broadcasts the signal fail, indicating that an exception has been thrown. All automata have a transition from every location (except for Terminated) to a location Timeout that is reached upon receiving the signal fail. For readability, these Timeout locations and all their incoming transitions are not shown in Figure 6 and Figure 7.
4.2. Abstractions
We have discussed the modelling of Java TCP/IP socket communications. We now discuss other aspects that have been abstracted away in the model. We note that the abstracted aspects are irrelevant for the analysis discussed in Section 6. In the model, the underlying orchestration automaton is abstracted together with the contracts of the services. Thus, all conditionals that are dependent from the underlying orchestration are abstracted as probabilistic choices. We assume that the orchestration has been correctly synthesised from the services contracts. This allows us to verify the interactions for any possible valid orchestration. If a specific orchestration would be modelled, then we would lose such generality. In particular, the payloads of the communications (e.g., which specific action, which choices) are abstracted. The conditions used to decide whether to perform a choice, an action or to stop are also abstracted away. The only modelled condition is that no two consecutive choices are allowed (i.e., after choosing, the chosen step must be performed). When executing a transition, the conditions used to check whether the label of the transition is an offer or a match are abstracted away in the model. Moreover, the identifiers of the services involved in performing a choice or an action (which are concretely extracted from the labels of the transitions) are also chosen non-deterministically. Indeed, all services are distinguished replicas of the same automaton. Finally, we model a single orchestration. In fact, when multiple orchestrations are executed, they operate independently of each other and can be verified individually.
4.3. Adequacy
Traceability and model-based testing have been used to ascertain the adequacy of the model with respect to the implementation. Traceability involves connecting the items of the abstract model with the lines of source code that are being abstractly represented. In [6] each transition of the model is traced back to the corresponding source code instructions using comments in the model. Traceability information is related to a specific version of the source code (i.e., b1a6954). For example, Figure 4 shows a portion of the automaton modelling the RunnableOrchestratedContract (discussed in details in the next section). Specifically, from state Ready a transition is executed which is traced back to line 98 of the class RunnableOrchestratedContract, also showed in Figure 4. Indeed, at line 98 the first read operation from the ObjectInputStream, reading from the TCP/IP socket, is performed by the service. This is abstracted in the model by the operation dequeue() of the automaton.
Figure 4. A fragment of the RunnableOrchestratedContract automaton equipped with traceability information (left) and a fragment of the class RunnableOrchestratedContract.java (right). The transition from state Ready is traced back to line 98 of the source code of the respective class.
Testing is used to show the adherence of the model to the actual implementation, and will be discussed in Section 7.
Remark
We highlight the advantages of employing graphical diagrams. When it comes to the implementation phase, developers typically work with the source code, which currently consists of 770 lines of code in the case of CARE. On the other hand, during the modelling phase with Uppaal, designers graphically edit automata. The automata depicted in Figure 6 and Figure 7 succinctly and accurately specify the interaction logic of CARE.
5. Formal model
The formal model of CARE is described. All models used in this paper together with the evaluated formulas are available in [6]. The network of automata is composed of a RunnableOrchestration automaton modeling the orchestrator, and each service is modeled by two automata: RunnableOrchestratedContract and SocketTimeout. Figure 5 shows the system declarations of the model. In particular, there is a single instance of RunnableOrchestration whilst there are N instance of RunnableOrchestratedContract and SocketTimeout. Indeed, both RunnableOrchestratedContract and SocketTimeout templates take as parameter id_t, which is a special global declaration of Uppaal (namely, typedef int[0,N-1] id_t) to instruct the model to produce N replicas of these template, where as stated above N is also a global constant of the model.
Figure 5. A snapshot of the Uppaal model showing the system declarations
Figure 6 displays the template automaton for the RunnableOrchestration (i.e., the orchestrator), while the template automaton for the RunnableOrchestratedContract (i.e., the service) has as parameter the id of the service (of type id_t) and is depicted in Figure 7. The behaviour according to the given configuration of action and choice is modelled inside each automaton. We anticipate that we will use variants of these two automata in Section 6, in order to perform different analyses. In particular, in Figures 6 and 7 the configurations of choice and action (for both the services and the orchestrator) are instantiated non-deterministically. Another version of the model is also available where the configuration options are also parameters of the templates. The repository [6] contains the model equipped with traceability information, i.e., each transition of the model has a comment describing the class and the lines of the source code that correspond to the specific behaviour of the transition. The models used for generating tests, together with the generated tests, are also available in [6].
The Uppaal model is composed of a list of global declarations, the two automata (with their local declarations) and the system set-up (i.e., the instantiation of the automata). Global declarations include the number of services N, the size of the buffers, the timeout threshold, the rates of the exponential distributions, two variables action and choice storing the corresponding configuration for all automata, and the communication buffers. Constants are also defined globally and their identifiers are in capital letters. Some names of locations are displayed in the automata for readability. Labels of transitions contain guards and effects and in a few cases also probabilistic selections.
5.1. Description of the automata
We now briefly describe Figure 6 and Figure 7. The initial state is depicted with a double circle. In the first transition of the orchestrator, one of four options encoding the combinations of choice and action is selected non-deterministically. The function initialize(conf) instantiates accordingly the choice and action global variables such that all have the same configuration, and also initialises the communication buffers.
From the location CheckCompatibility of the orchestrator a loop is executed to communicate with all services to check if all have the same configuration. The orchestrator sends the message ORC_CHECK and its configuration (action and choice). If a service has the same configuration an ACK is sent, or an ERROR message otherwise. In case of no errors, the orchestration starts. From the location Start the orchestrator internally decides (based on the orchestration) whether to perform a choice or an action. The choice of termination is modelled as a third alternative. As stated previously, after a choice is completed the orchestrator moves to a state where only an action or termination can be chosen. After an action is completed, the orchestrator returns to the Start location.
Figure 6. The RunnableOrchestration Uppaal template
Figure 7. The RunnableOrchestratedContract Uppaal template
In the case of termination, the orchestrator sends to all services an ORC_STOP message and the orchestration terminates. In the case of a choice, firstly all services receive the message ORC_CHOICE. If the choice is dictatorial, the orchestrator decides autonomously, and no further interactions are necessary. Otherwise, in the case of a majoritarian choice, the services involved in the choice are selected non-deterministically, and a message ORC_CHOICE is sent to them. Concretely, the involved services are those who perform an action in one of the outgoing transitions from the current state of the orchestration. The services that are not involved will receive a SKIP message. The involved services then receive from the orchestrator the available choices (i.e., the forward star of the current state). These concrete choices are abstracted by the constant CHOICES. Each involved service now replies with its choice, abstracted as a message SERVICE_CHOICE. After all involved services have voted, the orchestrator will decide accordingly.
In case of an action, the orchestration behaves differently depending on whether the configuration is centralised or distributed. In both cases, a probabilistic choice is made on whether the transition is an offer or a match. After that, the orchestrator picks non-deterministically an offerer and a requester (only in the case of a match transition) with consecutive identifiers (recall that their IDs are immaterial). In the case of a centralised offer action, the offerer receives from the orchestrator the invocation of the action, abstracted by the constant ACTION. At this point the offerer is still not aware of whether it is involved in an offer or a match transition. The reception of the NOPAYLOAD message (i.e., there is no payload from the requester) disambiguates the offerer who replies with a payload abstracted here by the constant OFFER.
In a centralised match the requester receives the action invocation, abstracted by the constant ACTION, and a SKIP command (i.e., the offer has not yet been generated). Note that in the implementation the service is informed of being an offerer or a requester upon receiving the action. Since the concrete action is abstracted, in the model this disambiguation occurs when receiving either the message SKIP (i.e., a requester) or the messages NOPAYLOAD (i.e., an offerer in an offer transition) or REQUEST (i.e., an offerer in a match transition). The requester replies with the message REQUEST (the concrete payload is abstracted away) that is forwarded by the orchestrator to the offerer, who receives in sequence the messages ACTION and REQUEST. Similarly to the previous case, the offerer sends to the orchestrator its offer payload (now based on the payload of the requester), abstracted again by the constant OFFER.
Concerning the distributed configuration, in case of an offer or match first the offerer receives the ACTION command from the orchestrator. In case of an offer, a TYPEOFFER message is received by the offerer followed by a NOPAYLOAD message. The offerer replies with an OFFER message. In case of a distributed match, the ACTION message is also sent to the requester and the TYPEMATCH message is sent to the offerer. The offerer opens a fresh port and communicates this port (abstracted by the constant PORT) to the orchestrator. The offerer waits for a connection from the requester. The orchestrator communicates to the requester (who was waiting after receiving the ACTION command) the address (constant ADDRESS) and PORT of the offerer. Now the requester and the offerer can interact without the orchestrator. The interactions between any two services in the orchestration all use two (one-position) buffers, implemented by the variables requester2offerer and offerer2requester. Indeed, it is never the case that some service interferes in a match in which it is not involved, and using different buffers for each pair of services would unnecessarily increase the state space. Initially, the requester sends its REQUEST payload to the offerer. The offerer replies to the requester with its payload OFFER (based on the request) and reaches location Done. The requester receives the payload and sends an ACK to the offerer and the orchestrator (who both terminate the execution of the match transition and returns to the Start and Ready state, respectively).
6. Analysis
We now describe the analyses we performed on the model. The modelling activity led to some issues in both the implementation and the model, which were all fixed (we remark that an already existing application was modelled). Other formal checks on the model were performed to ensure the properties described in this section (e.g., absence of deadlocks, absence of orphan messages). We have used Uppaal version 4.1.26-1, February 2022.
6.1. Validation through modelling
The first validation was performed during the modelling phase. Indeed, formal modelling requires an accurate analysis and review of the source code. Interactive simulation is used during modelling to animate and analyse the portion of the model designed so far, similarly to how the source code is debugged interactively (e.g., by choosing the next step). We note that in many model-based engineering tools, behavioural models (e.g., state charts) are validated by only relying on graphical interactive simulations [18]. Issues can be detected during this phase in particular if the source code has not been thoroughly tested, as was the case for CARE.
We report an issue detected in the source code during the modelling of the automaton in Figure 6, and more specifically during the modelling of the loop in which the orchestrator is reading the SERVICE_CHOICE messages sent by the involved services. In the implementation, the orchestrator was waiting for a choice from all services, also comprehending those who received a SKIP message. This means that a deadlock could occur in case there was a service not involved in a choice. Initially, this issue was undetected because the tests had all services involved in all choices. It was fixed thanks to the activity described in this paper.
On a side note, thorough testing is generally more time-consuming than designing a formal model similar to the one in this paper, and Uppaal has been used to automatically generate tests from the formal model.
6.2. Formal Verification
We discuss the formal verification performed on the model, encompassing both exhaustive and statistical model checking. In the initial phase, we employ a model variant that guarantees consistent configurations between the services and the orchestrator. In this variant, the selection of the choice and action configuration is performed non-deterministically by the first transition of RunnableOrchestration (variable conf, see Figure 6), and it is always consistent. Subsequently, we move to another variant where the configuration of choice and action are treated as parameters. In this case, we formally prove that when configurations do not match, an error state is reached.
Performances. Statistical model checking has been used to scale to larger systems, and the verification has been performed in a few seconds on a standard laptop. Conversely, exhaustive model checking necessitated more resources and has been limited to smaller configurations generating hundreds of millions of states (see below). In this case, the verification has been performed on a machine with Intel(R) Core(TM) i9-9900K CPU @ 3.60 GHz equipped with 32 GB of RAM. Uppaal has been configured for maximising the state space optimization and reusing the generated state space. Logs of the experiments are available in [6]. The summaries of the analysis performed with statistical (resp., standard) are in Table 1 (resp., Table 2).
Task Property Parameters’ tuning Interval Confidence
Buffer size Probability of filling a buffer queueSize set to 5 $[0,0.00996915]$ 0.95
Timeout Probability of timeout timeout set to 15 $[0,0.00996915]$ 0.95
Delays Probability of termination write and read set to 5 $[0.990031,1]$ 0.95
Orphan messages Probability of terminating with non-empty buffers - $[0,0.00999882]$ 0.995
Interference Probability of a service interfering on a match between two other services - $[0,0.00999882]$ 0.995
Table 1. A summary of the properties analysed with statistical model checking
Task Property Conf. Time Memory State space
Termination If orchestrator terminates then all terminate or a timeout occurs c1 3 m. 1.5 GB 52 M st.
c2 30 m. 12 GB 426 M st.
Deadlocks Absence of deadlocks (no enabled transitions, timeout, or termination) c1 3 m. 1.5 GB 25 M st.
c2 40 m. 13 GB 189 M st.
Orphan Messages Upon termination, no messages left in buffers c1 2 m. 1.5 GB 25 M st.
c2 27 m. 12.5 GB 189 M st.
Dummy Execution No dummy execution where no interaction occurs c1 3 sec. 1.4 GB 24 st.
c2 24 s. 10 GB 28 st.
Compatibility Check Mismatch leads to Error location Special config. 1 ms. 49 MB 79 st.
Table 2. A summary of the properties analysed with standard model checking that are satisfied by the model
6.3. Parameters Tuning
The verification process involves employing specific parameter configurations within the model. This encompasses various aspects, such as setting the delays in reading and writing, setting socket timeout thresholds, adjusting buffer sizes, assigning probability weights, and determining the number of services involved (i.e., the instantiations of the RunnableOrchestratedContract template). For deriving the desired set-up of parameters, we employ statistical model checking. If not stated otherwise, the parameters $\alpha$ and $\varepsilon$ of the statistical model checker are set to $0.05$ and $0.005$, respectively (see Section 3).
In this section, when reporting the evaluation of a probabilistic formula, instead of reporting an interval containing zero (e.g., $[0,0.00996915]$), sometimes, for conciseness, we will say that the property has a probability near zero. To be more precise, we mean that all runs do not satisfy the property, thus the probability lies within the interval $[0,0+\varepsilon]$ with probability $1-\alpha$. We remind readers that, unlike probabilistic model checking, in statistical model checking, the unknown probability value is estimated to lie within an interval with a given confidence.
It is crucial to note that we do not employ statistical model checking to determine values (such as buffers size) for use in the concrete implementation. Instead, these quantified values are used solely within the model. For instance, in the actual implementation, the size of Java TCP/IP socket buffers is fixed (more below). Our objective is to employ parameter configurations that ensure realistic modelling and improve the performances of the exhaustive model checking. Realistic modeling entails accurately representing the behavior of the real system, by reducing the probability of filling the buffers, timeouts, or excessive communication delays. Failure to maintain these conditions could potentially invalidate the results of the formal verification. Improving performances, on the other hand, entails reducing the state space of the model.
Firstly, the values of the probabilities weights pstop, pchoice, pnochoice, and paction (see Figure 6) can be tuned based on average values extracted from the orchestrations subject of the analysis. Indeed, as stated in Section 4, the underlying orchestration automaton is abstracted away. Note that, e.g., in location Start the probability of performing a choice is $\frac{\tt pchoice}{\tt pchoice+pnochoice}$ (pchoice is the probability of performing a choice and pnochoice is the probability of not performing a choice). For example, the orchestration in Figure 8 (right) can be modelled by tuning the probabilities to pstop=25, pchoice=1, pnochoice=0, and paction=75.
Figure 8. On the left side, the SocketTimeout automaton. On the right side, the orchestration automaton taken from the composition service example in [11]
Next, we address the buffer size (denoted as variable queueSize). It is important to note that the buffers in the model are represented by global arrays utilized by the automata for enqueuing and dequeuing values. These global arrays serve as models of the actual buffers found in Java TCP/IP sockets, where the default size is typically 8 KB. Our objective is to prevent unnecessary growth in the model’s state space while ensuring a low probability of the buffers filling up, similar to the behavior observed in the real application. The formula:
E[<=500; 10000](max: sum(i:int[0,N-1]) (sum(j:int[0,queueSize-1])(orc2services[i][j]!=nil)))
computes, using 10000 simulations of 500 time units, the expected maximum number of non-empty positions of the buffers used by the orchestrator to send messages to the services (called orc2services). For all statistical evaluations, we set the number of services to $10$. With buffer size set to $10$, this formula evaluates to $4.5\pm 0.019$. This indicates that, on average, the maximum number of utilized buffer positions is between 4 and 5. Consequently, in order to reduce the state space, it is safe to decrease the value of queueSize to less than 10 for the exhaustive model checking phase. This observation is further supported by the following formula:
Pr[<=500] (<>(exists (i:id_t) ror.isFull(i)))
which measures the probability that, within 500 time units, one of the arrays orc2services[i] becomes full (ror is the orchestrator automaton). In three separate experiments where queueSize was varied between 3, 4, and 5, this formula yielded the respective evaluation intervals of $[0.990031,1]$, $[0.00632077,0.0163207]$ (with $\alpha\!=\!0.005$), and $[0,0.00996915]$. Based on these results, if not stated otherwise, queueSize is set to 5 for the subsequent experiments.
Next, we consider the real-time behavior of the model and focus on determining appropriate values for the rates write and read, which represent the delays in writing to and reading from a buffer, respectively. The average message delay in Java TCP/IP sockets is affected by multiple factors, such as network conditions and server load. Delays can be sampled using tools like tcpdump, and the rate can be estimated by calculating the inverse of the sample mean. Additionally, we consider the variable timeout, which represents the timeout threshold. Our aim is to achieve three objectives. First, we strive to maintain a low probability of encountering timeouts. Second, we seek to ensure a high probability of terminating within a specific timeframe, which we have set to be 500 time units, corresponding to the duration used in our experimental setup. Third, we aim to keep the timeout threshold at a lower value in order to decrease the state space and facilitate model checking.
The formula:
Pr[<=500] (<> ror.Timeout)
measures the likelihood of one or more service sockets experiencing a timeout (recall that in this case a failure signal is broadcasted and all automata enter their respective Timeout location). The probability of all services and the orchestrator successfully concluding their operations is evaluated with the formula:
Pr[<=500](<>ror.Terminated&&(forall (i:id_t) ROC(i).Terminated))
(ROC is used as an abbreviation of RunnableOrchestratedContract). In different experiments where we varied the values of the pair (rate,timeout), specifically in $(5,14)$, $(4,15)$, $(5,15)$, respectively, the first formula (timeout probability) yielded the respective evaluation intervals $[3.06006e^{-06},0.00999494]$ (with $\alpha\!=\!0.005$), $[0.0433422,0.053341]$, and $[0,0.00996915]$, while the second formula (probability of termination) yielded $[0.990005,0.999997]$ (with $\alpha\!=\!0.005$), $[0.948625,0.958625]$ (with $\alpha\!=\!0.005$), and $[0.990031,1]$. Therefore, to fulfill the aforementioned three objectives, we have set the values of write and read to 5, while the value of timeout has been set to 15 for the subsequent experiments. Indeed, in the experiments we just mentioned, the value (5,15) for the pair (rate,timeout) yields the best values for both the probability of experiencing timeout and the probability of reaching termination.
Concerning the instances of the templates, in all experiments there is one instance of the orchestrator template and either 4 or 5 instances of the service template. For the exhaustive model checking phase, we used two small configurations of (number of services, buffers size). The first is c1=(4,5), the second is c2=(5,3). In fact, the configuration (5,4) remained inconclusive in the experiments due to the need for generating billions of states and the inadequate memory capacity of the utilized machine. The verification of larger configurations necessitates either relying exclusively on statistical model checking or employing more powerful machines.
6.4. Verification
Once the model’s parameter configuration is determined, our next step involves verifying additional formal properties.
Termination. We have already assessed the probability of non-termination and found it to be nearly zero. However, it may be worthwhile to conduct an exhaustive verification specifically for this property. The property that in all executions eventually all services and the orchestrator terminate is not valid. Indeed, as described in Section 4, the orchestration contract automaton is abstracted away and at each iteration, a choice is performed to decide whether to terminate or not. Hence, there exists an execution in which the orchestration never terminates. A milder property does hold:
ror.Stop-->((ror.Terminated&&(forall(i:id_t)ROC(i).Terminated))||
(exists(i:id_t)SocketTimeout(i).Timeout))
i.e., if the orchestrator decides to terminate then eventually all services and the orchestrator terminate. The formula p-->q is a shortcut for A[](p imply A<>q). Thus, this formula states that for all executions and for all states either the orchestrator is not in the location ror.Stop or all executions passing through that location will eventually lead to a state where all services and the orchestrator have terminated or a timeout failure is experienced. The formula holds in both configurations c1 and c2. The first (resp. second) configuration required roughly 3 (resp. 30) minutes, 1.5 (resp. 12) Giga of memory, and explored 52 (resp. 426) million states.
Absence of deadlocks. The likelihood of non-termination being very low also implies an almost negligible probability of encountering deadlocks. While it may be of interest to exhaustively prove the absence of deadlocks, the previous formula is insufficient for this purpose. Hence, to prove that there is no deadlock, we perform exhaustive model checking of the formula:
A[](not deadlock || (exists(i:id_t) SocketTimeout(i).Timeout) ||
(ror.Terminated && (forall (i:id_t) ROC(i).Terminated)))
the formula states that for all executions and for all states of the composed system, either there is always at least one enabled transition or either a timeout failure has been experienced or all services and the orchestrator are in the Terminated location. In the formula, not deadlock is a special predicate provided by Uppaal. As expected, this property is satisfied in both configurations. The first (resp. second) configuration required roughly 3 (resp. 40) minutes, 1.5 (resp. 13) Giga of memory, and explored 25 (resp. 189) million states. This also proves that in a correct configuration the Error location is never reached, because this would result in a deadlock and the above property would not be satisfied. This property is also satisfied when the size of the buffers is $3$. However, if we further reduce the size, then a deadlock occurs. This is because from location CheckCompatibility the orchestrator requires to insert three messages in the buffer of the receiver in one step. By dividing these three send operations in three non-committed transitions it is possible to further reduce the buffer size.
We report a modelling issue detected during model checking the above formula. In an earlier version, it was assumed that the socket mode was non-blocking (i.e., sending to a recipient with a full buffer would cause an error). This was modelled by making committed (C) all source states of transitions with sending operations. In this way, if the buffer of the receiver would not have enough free space then an attempt to send to the receiver would cause a deadlock. In fact, in this earlier version of the model the above formula (absence of deadlocks) was not satisfied, for any possible size of the buffer. The counterexample trace of the model checker helped to understand and eventually fix this issue. Basically, in the model configured with a majoritarian choice there exists a loop in which the orchestrator alternates choices and actions, and enqueues a sequence of ORC_CHOICE and SKIP messages to a service that never consumes them and is never involved in neither choices nor actions, thus eventually filling its buffer and deadlocking. A similar issue also exists in the case of a dictatorial choice.
Note that this kind of problems are hard to detect without model checking. Indeed, the counterexample trace was generated automatically, and the counterexample trace is composed of hundreds of steps. Without model checking this would require to manually execute each step of this trace, and the longer the trace the less chance to discover it. After we detected this issue, we analysed the underlying Java TCP/IP socket semantics [32] and fixed the model as described in Section 4 (i.e., by modelling these sockets with default blocking mode). Another fix could be to include an ack after the reception of a SKIP message (this would however require to modify also the implementation).
Absence of orphan messages. We now prove that upon termination of an orchestration no messages are left in any buffer, i.e., all messages are consumed. To expedite the verification process, we begin by conducting statistical model checking of the following property:
Pr[<=500](<>!allEmpty()&&ror.Terminated&&(forall(i:id_t)ROC(i).Terminated))
this property quantifies the probability of termination with at least one message remaining in any buffer. To verify whether all buffers are empty, we utilize the predicate allEmpty(). As expected, the probability is found to be nearly zero. We proceed with an exhaustive verification by employing the property
A[]((ror.Terminated && (forall (i:id_t) ROC(i).Terminated)) imply allEmpty())
the above formula can be read as follows: in all states of all executions either all buffers are empty or someone has not terminated yet. The above property is valid in both configurations, as expected. The first (resp. second) configuration required roughly 2 (resp. 27) minutes, 1.5 (resp. 12.5) Giga of memory, and explored 25 (resp. 189) million states. It is also possible to verify that there is no dummy execution in which the services and the orchestrator never interact (and thus all buffers are trivially empty). This can be verified with the formula:
E[] (allEmpty() && !ror.Timeout)
that, as stated above, checks if there exists an execution where all states have empty buffers (excluding the dummy execution scenario in which the timeout occurs at the beginning). As expected, this property is not valid in the model, for both configurations. The first (resp. second) configuration required roughly 3 (resp. 24) seconds, 1.4 (resp. 10) Giga of memory, and explored 24 (resp. 28) states.
No interference. When discussing the distributed match action in Section 5, we stated that it is never the case that some service interferes in a match in which it is not involved. This guarantees that it is safe to use two one-position buffers for all communications between any two services involved in a match. To verify this, we perform statistical model checking of the following formula
Pr[<=500](<>exists(i:id_t)(i<N-1&&(ROC(i).d1==TYPEMATCH||ROC(i).d1==ADDRESS||ROC(i).d1==PORT))
&&((ROC(i+1).d1==TYPEMATCH||ROC(i+1).d1==ADDRESS||ROC(i+1).d1==PORT))&&
(exists(j:id_t)(j!=i&&j!=i+1&&(ROC(j).d1==TYPEMATCH||ROC(j).d1==ADDRESS||ROC(j).d1==PORT))))
the formula measures the probability of reaching a state where one service (index j) is interfering on a match between two other services (indexes i and i+1). We recall that in the model two matching services have consecutive indexes. We detect a service to be involved in a distributed match when its temporary variable d1 has one of the three values (TYPEMATCH, ADDRESS, PORT). The probability is found to be nearly zero.
We also include in the repository [6] a version of the model where each pair of services has its own buffers. All results in this section also hold in that model.
Compatibility check. Next, we formally prove that if some service is not matching the configuration of the orchestrator, then the orchestration will not start and an Error location will always eventually be reached. Indeed, the possibility of mismatching configurations is allowed in the real system. However, in the model discussed in Section 5 this scenario is not possible because, by construction, all services and the orchestrator share the same configuration. The configuration is selected non-deterministically. In this way, for each formula being verified all possible configurations are checked automatically.
Only for this check the model has been slightly modified by adding two parameters action and choice to the templates and by updating accordingly the model. For this verification, the set-up of the system is of an orchestrator ror and three services alice, bob and carl and the size of each buffer is $4$:
ror = RunnableOrchestration(MAJORITARIAN_CHOICE,DISTRIBUTED_ACTION);
alice = RunnableOrchestratedContract(0,MAJORITARIAN_CHOICE,DISTRIBUTED_ACTION);
bob = RunnableOrchestratedContract(1,MAJORITARIAN_CHOICE,DISTRIBUTED_ACTION);
carl = RunnableOrchestratedContract(2,DICTATORIAL_CHOICE,DISTRIBUTED_ACTION);
ast = SocketTimeout(0);bst = SocketTimeout(1);cst = SocketTimeout(2);
system ror,alice,bob,carl,ast,bst,cst;
Note that, differently from Figure 5, the configurations of choice and action are now parameters assigned to each automaton. This allows us to assign a mismatching configuration to verify that the Error location will be reached. Indeed, in the above set-up carl has a different configuration. We use the formula:
A<>((ror.Error && carl.Error)||ror.Timeout)
stating that in all executions eventually the orchestrator and the service with a wrong configuration reach an Error location or a timeout is experienced. Alternatively, the formula:
A[](!ror.Start)
states that in all executions the Start location of the orchestrator is never traversed (i.e., the orchestration never starts). Both properties are satisfied in this setup, thus verifying the correctness of the compatibility check. The first (resp. second) formula required visiting 19 (resp. 79) states. Both formulae used roughly 48 Megabytes of memory and a few milliseconds of CPU. We have proved that in case of mismatching configurations, the orchestration will not start.
7. Testing
The model-based testing functionality of Uppaal has been employed to generate tests that demonstrate the model’s adherence to the actual implementation.
The abstract test cases generated by Uppaal have been transformed into concrete JUnit tests for the source code. Uppaal provides various methods for test generation (see Section 3). We utilized the generation from reachability queries (i.e., of the form E<>). In our models, these queries encode specific simulation traces that are relevant to the specific orchestration employed in the tests and are designed to cover all transitions of the model. The specific simulation traces also guide the generation of the concrete test cases from the abstract test cases. To achieve this, two global variables, namely steps[] and step, are introduced in the models to encode the desired orchestration in the query. The length of the array steps[] is equal to the length of the trace being simulated, and the variable step is a pointer to the current position in the array. Whenever a specific transition is executed, the variable steps[step] is updated with a value encoding the transition being executed, and the counter step is incremented. For example (see Figure 9), whenever the transition reaching state Error is executed, the value ERROR is assigned to steps[step]. As another example, whenever a centralised match transition is executed by the orchestration, the value CENTRALISED_MATCH is assigned to steps[step], encoding the execution of the request action in the match (see Section 5). In Section 7.3, we will provide an example showing how these additional variables are used in a query.
7.1. Testing the Orchestrator
Different variants of the model have been developed, each one equipped with specific test code to test one of the components. One version of the model is equipped with test code for testing the orchestrator (i.e., RunnableOrchestration). In this case, during concrete testing, the orchestrator will be the actual system, whilst the tester will mimic the services. Therefore, the test code will be generated from the services model. In this version, the automaton template RunnableOrchestratedContract contains test code in its transitions, whilst the automaton template RunnableOrchestration does not contain test code. However, useful comments are inserted as text code in some transitions of RunnableOrchestration to guide the generation of concrete tests from abstract tests. This mainly involves printing the actual value during simulation of variables offerer, requester and i used by the RunnableOrchestration as indices to identify the services the orchestrator is interacting with. In the generated tests, no timeout is experienced, therefore the SocketTimeoutAutomaton is a dummy automaton receiving signals from RunnableOrchestratedContract and doing nothing.
All transitions of RunnableOrchestratedContract containing a dequeue() operation in the effects will also contain as test code the instruction msg = (String) oin.readObject(). Basically, the abstract operation of reading a message from the queue is concretised as a read operation from the variable oin of class ObjectInputStream of the Java TCP/IP socket of the service. Similarly, transitions containing an enqueue effect will contain as test code instructions for writing into a socket. The information on the actual payload to be sent by the concrete services is abstracted away in the Uppaal model. For example, consider the transition reaching state Ready of RunnableOrchestratedContract with effect enqueue(SERVICE_CHOICE) (see Figure 9, top right corner). In this case, SERVICE_CHOICE is a constant abstracting from the choice that the service will be expressing, which will be based at runtime on the transitions outgoing the current state of the contract automaton of the service, abstracted in the Uppaal model. Accordingly, the test code generated from that transition is oout.writeObject(choice);oout.flush();. In this case, choice is a placeholder that needs to be replaced in the concrete test with the actual choice made at that point. Furthermore, oout is the ObjectOutputStream of the service’s TCP/IP socket, on which a write operation is performed.
Finally, some read operations and other transitions with guards will include assertions in their test code. These assertions will be concretised in the concrete tests by adding values that were abstracted away in the model and will be used by the JUnit tests to verify whether the tests are successful or not. For example, the transition of RunnableOrchestratedContract reaching state Terminated will contain in its test code the instruction assertEquals(msg,RunnableOrchestration.stop_msg);. In this case, it is tested whether the last received message is the constant identifying the termination of the current orchestration.
Figure 9. The RunnableOrchestratedContract Uppaal template used for testing the orchestrator
7.2. Testing the Services
For testing the RunnableOrchestratedContract services another version of the model is used. The orchestrator will be mocked by the test, whilst the services will be the actual classes of CARE. The abstract test code will be generated through the automaton RunnableOrchestration. The Uppaal automaton RunnableOrchestration will also be decorated with the effects on the array steps[step], similar to Figure 9. The generation of test code is similar to the case of testing the orchestrator, discussed previously. The main difference is that the orchestrator manages a list of input/output streams, one for each service. For example, when sending the message ORC_CHOICE to the services (see Figure 6), the transition with the effect enqueue(i,ORC_CHOICE) generates the test code oout.get(\$(ror.i)).writeObject(RunnableOrchestration.choice_msg); oout.get(\$(ror.i)).flush();, where the object oout is a list of ObjectOutputStream, one for each service. Similarly, the dequeue(i) effect generates the test code response = oin.get(\$(ror.i)).readObject();, where oin is a list of ObjectInputStream. Finally, similar to the case of testing the orchestrator, abstract values need to be instantiated in the concrete tests, and assertions are included in the tests to check whether the received messages are as expected.
For testing the interactions between two services in a distributed match, there are two further versions of the model. In these versions, the tester is represented by the orchestrator and one of the two interactive services, either the requester or offerer in the match.
In the following, we will provide an example on how the abstract test code is generated and transformed into a concrete test. We will focus on a specific version of the model, i.e., the one used for testing the orchestrator, and a specific set-up.
7.3. An example of test generation
We now discuss a query used for testing an orchestration whose set-up is dictatorial choice with centralised action. This means that the orchestrator will decide each choice and the orchestrator will act as a broker in a match between the two services. All versions of the model contains queries for testing the various configurations of the orchestrator (e.g., majoritarian choice, distributed action). We refer to [6] for more details about these other queries. In the concrete tests, the orchestrator will interact with two mocked services. Furthermore, the orchestration contract automaton, a parameter of the RunnableOrchestration object (see Figure 2), is dummy and is instrumental to cover the desired portions of code of CARE. In particular, the orchestration contract automato that will be used in the concrete tests accepts the trace (!euro,-)(?coffee,!coffee)[(!euro,-)]. The corresponding query used to generate the abstract test is:
E<>(alice.steps[0]==ORC_CHECK &&
bob.steps[1]==ORC_CHECK &&
alice.steps[2]==CENTRALISED_OFFER &&
alice.steps[3]==CENTRALISED_MATCH &&
bob.steps[4]==CENTRALISED_OFFER &&
alice.steps[5]==DICTATORIAL_CHOICE &&
bob.steps[6]==DICTATORIAL_CHOICE &&
alice.steps[7]==CENTRALISED_OFFER &&
alice.steps[8]==DICTATORIAL_CHOICE &&
bob.steps[9]==DICTATORIAL_CHOICE &&
alice.steps[10]==ORC_STOP &&
bob.steps[11]==ORC_STOP)
The query above simulates a run accepted by the orchestration contract automaton. We now analyse the query. As discussed in Section 5, the orchestration always starts by checking if the configurations of the services and the orchestrator are matching. Therefore, both services are instructed to check their configuration (alice.steps[0]==ORC_CHECK and bob.steps[1]==ORC_CHECK). In this test, the configuration is of centralised action and dictatorial choice. After the configuration checks, the orchestration starts. In the query, Alice is instructed to execute the offer action !euro with alice.steps[2]==CENTRALISED_OFFER. This corresponds to execution of the offer label (!euro,-) in the orchestration (Alice is the first service). After that, the match label (?coffee,!coffee) of the orchestration is executed. The match starts with Alice performing the request ?coffee, encoded in the query as alice.steps[3]==CENTRALISED_MATCH. After Alice performs the request action, it is Bob’s turn to match the request of Alice with the corresponding offer action!coffee. In Figure 9, it is possible to observe that the transition of the RunnableOrchestratedContract Uppaal template covering the execution of an offer action in both the cases of offer label or match label (respectively, d1==NOPAYLOAD or d1==REQUEST, see Section 5) has in its effect the instruction steps[step]=CENTRALISED_OFFER. Therefore, Bob is instructed to execute the offer action !coffee in the query with bob.steps[4]==CENTRALISED_OFFER.
After the offer label (!euro,-) and the match label (?coffee,!coffee) have both been executed, a choice has to be made on whether to execute the offer label (!euro,-) or terminate (recall that the trace (!euro,-)(?coffee,!coffee)[(!euro,-)] is accepted by the orchestration contract automaton). In this set-up (i.e., dictatorial choice), the orchestrator performs the choice and communicates it to the services. The orchestrator will decide the first time to execute the offer label (!euro,-) and the second time to terminate. thus covering both cases. The first choice is encoded in the query with alice.steps[5]==DICTATORIAL_CHOICE and bob.steps[6]==DICTATORIAL_CHOICE. As before, the offer label is encoded in the query with alice.steps[7]==CENTRALISED_OFFER. Finally, the orchestrator performs the choice to terminate. This is encoded in the query with alice.steps[8]==DICTATORIAL_CHOICE and bob.steps[9]==DICTATORIAL_CHOICE (the orchestrator communicates the choice to the services) and alice.steps[10]==ORC_STOP, bob.steps[11]==ORC_STOP (both services terminate their execution).
We now discuss the abstract test code generated by the query and how it is concretised into a JUnit test for the real system. Uppaal also allows the insertion of abstract test code in the system declaration of a model, which will be printed out in the abstract test either at the start (i.e., pre-fix) or upon termination of the test (i.e., post-fix). In all tests, pre-fix test code is used to declare some variables that will be utilised during the test. The first transitions executed during the test are those related to the configuration check. Both Alice and Bob generate the same portion of abstract test code, reported below.
{minted}
[mathescape,fontsize=,linenos]java msg = (String) oin.readObject();
//INIT CHECK assertEquals(msg,RunnableOrchestration.check_msg);
String received_choiceType = (String) oin.readObject(); String received_actionType = (String) oin.readObject(); assertEquals(received_actionType,actionType); assertEquals(received_choiceType,choiceType);
oout.writeObject(RunnableOrchestration.ack_msg); oout.flush(); //STOP CHECK
In Figure 9, for both services, after the transition from the initial state Ready is executed, the transition with the guard d1==ORC_CHECK is fired, followed by the subsequent transition, and finally, the transition having the effect steps[step]==ORC_CHECK is executed. In order to satisfy the query these four transitions must first be executed by Alice (when step=0) and then by Bob (when step=1).
The first transition performs a dequeue operation, which appends the readObject instruction (line 1). The second transition executed, with the guard d1==ORC_CHECK, generates an assertion to check whether the message received by the orchestrator is the command for performing the configuration check (line 4). The next transition performs two dequeue operations, generating two readObject instructions, followed by two assertions to check whether the two messages are those expected by the orchestrator (lines 6-9). Recall that the orchestrator communicates the choice and action configuration (see Figure 6) to all services. It is important to note here that the actual values to check (actionType and choiceType) will be instantiated in the concrete test, discussed next. The fourth executed transition performs an enqueue operation, generating the test code writeObject, which writes the corresponding acknowledgement message into the socket (lines 11-13).
We now move to the structure of the concrete JUnit test and the concretisation of the abstract test code discussed until now. After that, we will discuss the remaining part of the abstract test code and its concretisation. The concrete test file is available at [6] and is called DictatorialCentralisedRunnableOrchestrationTest.java. All concrete tests used for testing the orchestrator follow the same schema, reported below (for the case of dictatorial choice and centralised action).
{minted}
[mathescape,fontsize=,linenos]java @Test public void test() throws IOException, ClassNotFoundException, InterruptedException alice = new Thread(()-¿ check(8080); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8080)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_alice(oin,oout);
catch (IOException — ClassNotFoundException e) throw new RuntimeException(e); );
bob = new Thread(()-¿ check(8081); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8081)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_bob(oin,oout); catch (IOException — ClassNotFoundException e) throw new RuntimeException(e); ); alice.start(); bob.start(); RunnableOrchestration ror = new DictatorialChoiceRunnableOrchestration(new Agreement(), adc.importMSCA(dir+”Orchestration.data”), Arrays.asList(null,null), Arrays.asList(8080,8081), new CentralisedOrchestratorAction());
Thread tror = new Thread(ror); tror.start(); tror.join();
The test is organised as follows. Both Alice and Bob services are encoded as parallel threads. Alice’s thread is declared at lines 3–16, whilst Bob’s thread is declared at lines 18–30. They are both executed at lines 31–32. After that, the orchestration is declared at lines 33–37 and executed as another parallel thread (lines 39–41). It is important to note that, whilst Alice and Bob are two mocked services, the orchestrator is the real RunnableOrchestration class of CARE. In this case, at lines 33–37, we can see that the RunnableOrchestration object ror is instantiated with dictatorial choice and centralised action. The other tests for the orchestrator will have different instantiations (majoritarian choice and/or distributed action). Furthermore, it uses the orchestration contract automaton discussed above, stored in the file Orchestration.data. The ror object also takes as a parameter the ports of the services. Alice is listening at port 8080, and Bob is listening at port 8081.
Alice’s thread starts by invoking the method check, passing its port as an argument (line 4). After that, the socket and object streams are initialised, and the method uppaal_alice is invoked, passing the input and output object streams as arguments. Bob’s thread is equivalent. The method check (not reported for brevity), similarly to lines 5–10 of method test above, opens a socket awaiting connection from the orchestrator and invokes the method uppaal_check. The method uppaal_check contains the concretisation of the abstract test code discussed above, which is used for checking the configuration of the corresponding service. The methods uppaal_alice and uppaal_bob (line 11 and line 26) contain the remaining concretisation of the abstract test code generated by Alice and Bob, respectively, and will be discussed in the following. The method uppaal_check is reported below.
{minted}
[mathescape,fontsize=,linenos]java private void uppaal_check( ObjectInputStream oin, ObjectOutputStream oout) throws IOException String choiceType = ”Dictatorial”; String actionType = ”Centralised”; String msg; msg = (String) oin.readObject(); assertEquals(msg,RunnableOrchestration.check_msg); String received_choiceType = (String) oin.readObject(); String received_actionType = (String) oin.readObject(); assertEquals(received_actionType,actionType); assertEquals(received_choiceType,choiceType); oout.writeObject(RunnableOrchestration.ack_msg); oout.flush();
It is easy to see that this method includes the abstract test case instructions discussed above. In this case, the concretisation of the abstract test case simply instantiates the variables choiceType and actionType to their respective values (lines 2-3). Note that the method uppaal_check will be invoked by both services.
We have discussed the main structure of the concrete test, the abstract test case instructions covering the configuration check and their concretisation. We now discuss the remaining instructions of the abstract test case and their concretisation in the methods uppaal_alice and uppaal_bob.
After the orchestration check has been performed, the orchestration starts, and the offer label (!euro,-) of the orchestration will be executed. The following instructions are appended to the abstract test case being generated.
{minted}
[mathescape,fontsize=,linenos]java //centralised offerer 0
msg = (String) oin.readObject();
assertEquals(msg, expected action);
msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,expectedPayload);
oout.writeObject(INSERT OFFER); oout.flush();
The transition outgoing state CentralisedOffer of the RunnableOrchestration automaton (see Figure 6) has as test code //centralised offerer \$(ror.offerer). This will append the comment at line 1 in the abstract test, indicating that the next instructions of the abstract test will cover the first offer action performed by Alice (service at index 00). After performing the check, both Alice and Bob are again in state Ready (Figure 9). The first readObject operation (line 3) is again appended by the outgoing transition from state Ready of Alice. The next transition executed by Alice, with guard d1==ACTION, will append the instruction at line 5. The assertion will check whether the message received by the orchestrator will be the expected action, which will be concretised in the concrete test with the value "euro" (i.e., the orchestrator is invoking the method euro of Alice).
After that, Alice automaton will execute the centralised action, and the subsequent transition will perform a dequeue operation, appending to the abstract test case the instructions at lines 7-8. In the assertion, expectedPayload is a placeholder that will be concretised with value null. Recall that in CARE each action of the contract automaton of each service is in correspondence with a method of the corresponding service class (see Section 3), which has input parameters and a returned value. In this case, the parameters and returned values of the method in correspondence with the offer euro of Alice are dummy (i.e., concretised with null values). The next transition with effect enqueue is executed by Alice. This transition also has in the effect the instruction steps[step]=CENTRALISED_OFFER, thus fulfilling the query (at this point of the simulation it holds step=2). The execution of the transition will append instructions at lines 10-11 to the abstract test case. Indeed, the enqueue operation corresponds to a writeObject operation in the test. The placeholder INSERT OFFER will be replaced by the concrete value null in the concrete test.
This concludes the test code for the first offer label (!euro,-) of the orchestration. Next, the match label (?coffee,!coffee) is executed. The instructions appended to the abstract test case during the execution of the match label are reported below.
{minted}
[mathescape,fontsize=,linenos]java //requester: 0
msg = (String) oin.readObject();
assertEquals(msg, expected action);
msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,expectedPayload);
oout.writeObject(INSERT REQUEST); oout.flush();
//matching offerer: 1
msg = (String) oin.readObject();
assertEquals(msg, expected action);
msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,expectedPayload);
oout.writeObject(INSERT OFFER); oout.flush();
//requester: 0
msg = (String) oin.readObject(); //receiving the matching offer assertEquals(msg, expected payload);
The three comments (lines 1,13,25) are appended by the RunnableOrchestration automaton, after state CentralisedMatch is reached (see Figure 6), and are useful to identify the instructions generated by Alice (requester 0) and Bob (matching offerer 1). The instructions at lines 3-11 are appended by Alice when executing the request action ?coffee. The placeholder expectedPayload at line 8 will be concretised with the value "coffee" (i.e., the orchestrator is invoking the method coffee of Alice), whilst the placeholder at line 10 will be replaced with the concrete value "request payload" (i.e., the request sent by Alice to Bob is the dummy String "request payload"). Indeed, although the services are mocked by the test, in case of a match these values will be handled by the real CARE orchestrator, which will act as a broker between the two services. After executing the transition with effect steps[step]=CENTRALISED_MATCH (Figure 9), thus fulfilling the query, Alice waits for Bob to send the offer. The instructions at lines 15-23 are appended by Bob when executing the offer action !coffee. Note that these are identical to the previous offer action !euro fired by Alice, because the same transitions of Figure 9 have been executed. The placeholders at line 17, line 20 and line 22 will be replaced by the concrete values "coffee" (the orchestrator invokes the method coffee of Bob), "request payload" (the payload sent by Alice to Bob is checked) and "offer payload" (the dummy reply of Bob to Alice), respectively. After that, the instructions at lines 27-28 are executed by Alice for completing the request action. The placeholder at line 28 will be replaced by the concrete value "offer payload" (the reply of Bob to Alice is checked).
This concludes the match (?coffee,!coffee). As stated earlier, the orchestrator will now perform a choice on whether to continue and execute the offer label (!euro,-) or terminate. For brevity, we only report the abstract test code dumped by a service when executing a dictatorial choice, reported below.
{minted}
[mathescape,fontsize=,linenos]java msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.choice_msg);
The instruction at line 1 is again appended when executing the dequeue effect outgoing state Ready (Figure 9). After that, the transition with guard d1 == ORC_CHOICE is executed, which will append the test code at line 3. The next executed transition, with effect steps[step]=DICTATORIAL_CHOICE, has no test code generated but it is used to fulfill the query. No other test code is generated for the dictatorial choice, as the orchestrator decides on its own the next move. This test code is generated by both Alice and Bob, performing in turn the choice. In this example, after the first dictatorial choice, the orchestrator will execute again the offer label (!euro,-), and again a dictatorial choice. We have already covered the test code that will be generated. After that, the orchestrator decides to terminate. The test code generated by each service is:
{minted}
[mathescape,fontsize=,linenos]java msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.stop_msg);
The instruction at line 1 is the first dequeue operation as above. The instruction at line 3 is instead appended by the test code of the transition with guard d1==ORC_STOP (Figure 9). This transition has also in its effect the instruction steps[step]=ORC_STOP, thus fulfilling the query. Note that both values RunnableOrchestration.choice_msg and RunnableOrchestration.stop_msg are constants of the class RunnableOrchestration of CARE, used to identify the two commands of the orchestrator.
We have already discussed how the values of the abstract test code are concretised. For completeness, we report below the code of the method uppaal_alice discussed before, which contains test code for the service Alice, obtained by concretising the abstract test code as discussed above.
{minted}
[mathescape,fontsize=,linenos]java private void uppaal_alice( ObjectInputStream oin, ObjectOutputStream oout) throws IOException String msg; msg = (String) oin.readObject(); assertEquals(msg, ”euro”); msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,null); oout.writeObject(null); oout.flush();
msg = (String) oin.readObject(); assertEquals(msg, ”coffee”); msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,null); oout.writeObject(”request payload”); oout.flush();
msg = (String) oin.readObject(); //receiving the matching offer assertEquals(msg, ”offer payload”);
msg = (String) oin.readObject(); assertEquals(msg,RunnableOrchestration.choice_msg);
msg = (String) oin.readObject();
//if the orchestrator decides not to terminate while (msg.equals(”euro”)) msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg, null); oout.writeObject(null); oout.flush(); msg = (String) oin.readObject(); assertEquals(msg, RunnableOrchestration.choice_msg); msg = (String) oin.readObject();
assertEquals(msg,RunnableOrchestration.stop_msg);
Note that the above code allows Alice to perform more than a single offer "euro" before terminating (lines 26 34). In this way it is possible to handle all possible traces of the orchestration automaton, and all internal choices of the orchestrator.
Similarly, we report below the code of the method uppaal_bob.
{minted}
[mathescape,fontsize=,linenos]java private void uppaal_bob(ObjectInputStream oin, ObjectOutputStream oout) throws IOException String msg;
msg = (String) oin.readObject(); assertEquals(msg, ”coffee”); msg = (String) oin.readObject(); // reading payload centralised action assertEquals(msg,”request payload”); oout.writeObject(”offer payload”); oout.flush();
msg = (String) oin.readObject(); assertEquals(msg,RunnableOrchestration.choice_msg);
//in case the orchestrator decides not to terminate do msg = (String) oin.readObject(); while (msg.equals(RunnableOrchestration.choice_msg));
assertEquals(msg,RunnableOrchestration.stop_msg);
Similar to Alice, also the code of Bob allows to handle many unfoldings of the final loop of the orchestration.
This concludes the example of test generation. Recall that this is one query of one version of the model (for testing the orchestrator). There are other queries used for testing the other configurations (e.g., majoritarian choice, distributed offer) and other versions of the model for testing the other components. However, the generation of abstract tests and their concretisation follows the same approach covered by this example. All these versions and the concrete tests are available at [6].
7.4. Coverage
The coverage information generated by IntelliJ is in Figure 10.
Figure 10. The coverage information generated by IntelliJ
The generated abstract tests cover all transitions of the model and all interactions between the orchestrator and the services. The code coverage indicates that the concrete tests derived from the model cover a significant portion of the source code. This suggests that the model is not excessively abstract compared to the actual implementation.
On a side note, the modelling and verification allowed us to fix some issues (see Section 6). This would not have been possible in case the abstraction were too coarse.
8. Conclusion and Future Work
We have presented the formal modeling and verification of the Contract Automata Runtime Environment (CARE), an open-source platform. Both statistical and exhaustive model checking techniques have played a crucial role in formally verifying numerous desired properties of the modeled system, such as the absence of deadlocks, while also enhancing the accuracy of the formal model. Statistical model checking has been employed to fine-tune parameter settings within the formal model, such as the buffer size. The adequacy of the abstract model has been described, and the transitions of the formal model have been linked to the corresponding lines of source code. The tests generated from the formal model have been employed to test the source code.
At present, the different artifacts such as the model, source code, tests, and tracing information are manually kept aligned. This process demands substantial effort whenever a new version of CARE is introduced, as each artifact needs to be updated accordingly. Future work involves studying techniques for automatic alignment of these artifacts. Furthermore, it would be interesting to exploit the facilities provided by the recently introduced tool Uppex [43] to factorize the different variants of the model discussed in the paper under a single, configurable Uppex model, and to automate the selection of a configuration.
Acknowledgment
Part of this study was carried out within the MUR PRIN 2022 PNRR P2022A492B project ADVENTURE (ADVancEd iNtegraTed evalUation of Railway systEms) funded by the European Union - NextGenerationEU, and the CNR project “Formal Methods in Software Engineering 2.0”, CUP B53C24000720005.